# One-Step Model Predictive Path Integral for Manipulator Motion Planning Using Configuration Space Distance Fields

### v1

@@ IV Numerical Simulations
The results are summarized in Table LABEL:tab:exp2. Our method achieves consistently high success rates. The path length results indicate that, for the QP method, longer trajectories are more likely to be trapped in local optima. In terms of computational efficiency, our approach achieves an average of 776 Hz in Scenario 1 (with 60 obstacles) and 718 Hz in Scenario 2 (with 123 obstacles).
We further compare the computation frequency with the MPPI method [15], which is similar to [13]. Since the original MPPI framework is tailored for end-effector tracking, its cost function is not directly applicable to joint-to-joint navigation. Therefore, we do not report success rates or joint-space path lengths. Instead, we evaluate the end-effector positions corresponding to target joint configurations and use them for planning. The results show that MPPI achieves only 61 Hz on average in Scenario 1 and 28 Hz in Scenario 2. As the number of obstacles increases, collision-checking time also grows. In contrast, [15] reported that MPPI can reach up to 500 Hz on an RTX 4090 when the number of obstacles is relatively small. Compared to this, our method sustains significantly higher control frequencies while maintaining near-perfect success rates.
@@ V Conclusion
We have proposed a method that integrates CDFs with MPPI. Our approach effectively leverages the complementary strengths of both CDFs and MPPI.
The proposed method achieves control frequencies exceeding 750 Hz with consistently high success rates, significantly outperforming the QP baseline, which is prone to local optima. Compared to the standard MPPI framework, our approach simplifies the cost function, unifies its units, and exploits the gradient information provided by CDFs to shorten the prediction horizon, thereby attaining substantially faster control frequencies.
Future work will focus on further optimizing the CDF representation. One direction is to extend CDFs beyond point-based sampling in the workspace. In our current experiments, we approximate obstacles using the centers of bounding spheres; a promising extension would allow CDFs to compute the shortest distance in configuration space to entire spherical regions. Another direction is to improve robustness near obstacles. When the CDF values are small—i.e., when the manipulator is close to obstacles—collisions can occur due to the highly non-convex nature of the configuration space. Developing strategies that enable safe initialization and reliable avoidance in such regions is an important avenue for future research.
@@ VI Appendix
$-\nabla_{q}f_{c}(\mathbf{p},\mathbf{q})\,\mathbf{u}_{k}\Delta t$ $\leq\ln\!\big{(}f_{c}(\mathbf{p},\mathbf{q})+\gamma\big{)},$ (22)
where $q_{k}$ and $u_{k}$ denote the state and control input at time step $k$, $H$ and $R$ are positive definite weighting matrices for tracking error and control effort, $e(\mathbf{q}_{k})=\mathbf{q}_{k}-\mathbf{q}_{\text{desire}}$ is the configuration error, $f_{c}(\mathbf{p},\mathbf{q})$ is the CDF, $\Delta t$ is the time step, $\gamma$ is a safety margin, and $\mathcal{Q},\mathcal{U}$ are admissible sets of state and control.
@@
$c_{\text{goal}}$ $=\lVert\mathbf{q}_{H}-\mathbf{q}_{\text{goal}}\rVert_{2},$ (23)
@@
$c_{\text{joint}}$ $=\sum_{h=1}^{H}\lVert\max(0,\mathbf{q}_{\min}-q_{h})+\max(0,\mathbf{q}_{h}-\mathbf{q}_{\max})\rVert_{2}^{2},$ (25)
$c_{\text{stay}}$ $=\frac{1}{\lVert\mathbf{q}_{H}-\mathbf{q}_{0}\rVert_{2}+\varepsilon},$ (26)
@@ VI-A Hyperparameters
$\alpha_{\mu}$ $0.3$
$\alpha_{\Sigma}$ $0.8$
Ours (Algorithm 1) $\alpha_{1}$ $2000$
$\alpha_{2}$ $1000$
@@
$\alpha_{\mu}$ $0.5$
$\alpha_{\Sigma}$ $0.5$
@@
QP $H$ $\mathrm{diag}(150,\,190,\,80,\,70,\,70,\,90,\,100)$
@@
Ours (Algorithm 1) $\alpha_{1}$ $2000$
$\alpha_{2}$ $1000$
@@
$\alpha_{\mu}$ $0.5$
$\alpha_{\Sigma}$ $0.5$

### v2

@@ V Conclusion
We have proposed a method that integrates CDF with MPPI. Our approach effectively leverages the complementary strengths of both CDF and MPPI.
The proposed method achieves control frequencies exceeding 750 Hz with consistently high success rates, significantly outperforming the QP baseline, which is prone to local optima. Compared to the standard MPPI framework, our approach simplifies the cost function, unifies its units, and exploits the gradient information provided by CDF to shorten the prediction horizon, thereby attaining substantially faster control frequencies.
Future work will focus on further optimizing the CDF representation. One direction is to extend CDF beyond point-based sampling in the workspace. In our current experiments, we approximate obstacles using the centers of bounding spheres; a promising extension would allow CDF to compute the shortest distance in configuration space to entire spherical regions. Another direction is to improve robustness near obstacles. When the CDF values are small—i.e., when the manipulator is close to obstacles—collisions can occur due to the highly non-convex nature of the configuration space. Developing strategies that enable safe initialization and reliable avoidance in such regions is an important avenue for future research.
@@ VI Appendix
$-\nabla_{q}f_{c}(\mathbf{p},\mathbf{q})\,\mathbf{u}_{k}\Delta t$ $\leq\ln\!\big(f_{c}(\mathbf{p},\mathbf{q})+\gamma\big).$ (22)
where $\mathbf{q}_{k}$ and $\mathbf{u}_{k}$ denote the state and control input at time step $k$, $H$ and $R$ are positive definite weighting matrices for tracking error and control effort, $e(\mathbf{q}_{k})=\mathbf{q}_{k}-\mathbf{q}_{f}$ is the configuration error, $f_{c}(\mathbf{p},\mathbf{q})$ is the CDF, $\Delta t$ is the time step, $\gamma$ is a safety margin, and $\mathcal{Q},\mathcal{U}$ are admissible sets of state and control.
@@
$c_{\text{goal}}$ $=\lVert\mathbf{q}_{H}-\mathbf{q}_{f}\rVert_{2},$ (23)
@@
$c_{\text{joint}}$ $=\sum_{h=1}^{H}\lVert\max(0,\mathbf{q}_{\min}-\mathbf{q}_{h})+\max(0,\mathbf{q}_{h}-\mathbf{q}_{\max})\rVert_{2}^{2},$ (25)
$c_{\text{stay}}$ $=\frac{1}{\lVert\mathbf{q}_{H}-\mathbf{q}_{0}\rVert_{2}+\varepsilon}.$ (26)
@@ VI-A Hyperparameters
$\alpha_{\boldsymbol{\mu}}$ $0.5$
$\alpha_{\boldsymbol{\Sigma}}$ $0.3$
Ours (Algorithm 1) $\alpha_{1}$ 2020
$\alpha_{2}$ $10$
@@
$\alpha_{\boldsymbol{\mu}}$ $0.5$
$\alpha_{\boldsymbol{\Sigma}}$ $0.5$
@@
QP $H$ $\begin{aligned} &\mathrm{diag}(150,\,190,\,80,\,70,\\ &70,\,90,\,100)\end{aligned}$
@@
Ours (Algorithm 1) $\alpha_{1}$ 2020
$\alpha_{2}$ $10$
@@
$\alpha_{\boldsymbol{\mu}}$ $0.5$
$\alpha_{\boldsymbol{\Sigma}}$ $0.5$